# ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction

### v1

@@ 7 Mass accretion rates in the V-shaped structure
Based on the idea that the V-shapes are a result of gas flowing toward cores, in this section we provide estimates of their mass accretion rates () for NH and H.
7.1 NH mass accretion rate
To derive the associated to the main blue V-shape (Fig. 10), we need to estimate its NH mass. For this procedure we use PySpecKit with the n2hp_vtau fitter, to fit the full NH line. The fitted parameters (see below) allow us to derive the (NH). The V-shape structure contains 77, 143, and 25 NH spectra with one, two, and three velocity components respectively. The bluest velocities in the three velocity component spectra accounts for 2% of the total number of velocities in this V-shape. For this reason, we model the full NH hyperfine line structure with one and two velocity components. In Table 2 we list the parameters and ranges used for this procedure.
After obtaining the modeled NH cube, we remove modeled components where , where and represent the estimated opacity and its associated error, respectively. This criterion is to ensure that we use reliable fitted parameters to determine our (NH) values. For the fitting of two NH components we only analyze the most blueshifted component. The resulting opacities follow a log-normal distribution with a peak at .
We derive the (NH) of the V-shape by using Eq. 5 (Caselli et al., 2002b):
Where , T, and are the opacity, excitation temperature, and velocity dispersion respectively, obtained from the full line fitting. The Planck and Boltzmann constants are represented by and respectively, and are the frequency and wavelength of NH, is the Einstein coefficient of the NH (10) transition, and are the statistical weights of the lower and upper energy levels, is the partition function, and is the energy of the lower level.
From the above procedure, inside the V-shaped structure, we get a total (NH) = 5.2410 cm and a total (NH) = 5.910 M. We use the average timescale of the VGs from § 5.3 (Fig. 10), = 64.5 kyr, to determine the (NH) as:
Note that the (NH) estimate (and (H) below) should be multiplied by , in order to account for the unknown inclination angle () of the protocluster/filaments relative to the POS.
Excitation temperature [K] | (T) | 2.73 80 |
Opacity | () | 0.01 40 |
Centroid velocity [km s] | () | 25 15 |
Velocity dispersion [km s] | () | 0.20 3 |
7.2 H mass accretion rate
@@
To ensure that we estimate the V-shape (H) on the same area as in the NH hyperfine line fitting, we use the CASA task imregrid to obtain the continuum-derived (H) map from Díaz-González et al. (2023) at the resolution of the NH data. We determine that in this V-shape the total (H 1.03 cm. Here, we derive a (H) map using Eq. 7:
@@
where from this (H) map we consider only the points that are part of the V-shape. To determine the mass associated to flowing motions we subtract the core masses (from Louvet et al. submitted) that are located inside this V-shape. Note here that this mass map is an upper limit given that we do not apply a background correction. We obtain a total of (H 54 M. Considering used in Eq. 6, we derive the (H) as:
@@
We use the procedure described in this section to estimate the (H) of other eight V-shapes shown in Fig. D.2. We include these values in Table D.1. The average (H) of these V-shapes is 3.4 10 M yr. We estimate the free-fall time of all 45 1.3 mm cores from Louvet et al. (submitted). These values present large scattering, ranging from M yr, with 28 of these cores presenting (H) M yr. For cores 2 & 3, the average (H) is 15.5 M yr, about twice the (H) of the main blue V-shape, located between these two cores.
7.3 NH relative abundance
@@
For the main blue V-shape (Fig. D), we derive the NH relative abundance (NH), using the NH and H column densities obtained above, as:
The NH) value obtained above is consistent, within the scatter, with other estimates in massive Galactic star formation regions in the ranges of (1.6 - 3.8) 10 (Caselli et al., 2002a; Henshaw et al., 2014, Sandoval-Garrido et al. in prep.).
@@ 8 Discussion
The V-shaped velocity gradients described in this work have been detected across multiple Galactic star forming system. Stutz & Gould (2016) introduced the Slingshot mechanism in the Integral Shaped Filament (ISF) located in Orion A. They show undulations of the region in both position and velocity, suggesting that these features appear to be ejecting protostars (see their Fig. 2). Furthermore, Stutz (2018) characterize a standing wave in the neighborhood of the ISF, consistent with the Slingshot mechanism. It is possible that the undulations in the works above might result in the observed V-shaped structures seen in different studies (see below). González Lobos & Stutz (2019) identify six evenly spaced (every 0.44 pc) velocity peaks along the spine of the ISF in Orion A. They suggest that this periodicity is consistent with the wave-like perturbation in the gas caused by the Slingshot mechanism. In Álvarez-Gutiérrez et al. (2021) they analyze the L1482 filament located in the California Molecular Cloud. In all of the analyzed tracers there is a clear velocity peak in their north region (length 1.8 pc, mass M). This subregion contains a higher gas density and higher number of YSOs compared to the more quiescent south part.
@@
While the two regions described above are considered nearby (both at D 500 pc), more distant regions also present these velocity features which we list below. Zhou et al. (2022) study the velocity profiles along filaments from the ATOMS survey (Liu et al., 2020b). The median mass of their sources is 1.4 M with a median length of the filaments at pc. By analyzing the HCO (10) emission they find converging VGs along filaments (see their Figs. 6 & 10), which they also detected using simulations from Gómez & Vázquez-Semadeni (2014). These VGs at scales comparables to the V-shapes presented here are consistent with our VGs estimates (see their Fig. 7 & 8). In Zhou et al. (2023) they analyzed CO (21) APEX/LAsMA data of the G333 complex. They identify multiple V-shaped VG (see their Fig. 7) which they describe as the PV projection of a funneling structure in PPV space (see their Fig. 9). The origin of this structure is due to material inflowing towards the central hub and also due to gravitational contraction of star-forming clouds or clumps.
@@
Redaelli et al. (2022) use ALMA NH (10) isolated component data of the high-mass (5200 M) clump AGAL014.492-00.139 identifying multiple coherent structures in PPV space (trees “B” and “G”; right panel of their Fig. 7 & 9). These are characterized by multiple undulations, and possible V-shaped VGs. For their “G” PV distribution, they suggest that one scenario is where the dense gas is flowing along the filament (of length 0.2 pc) from protostar “p3” towards the protostar “p2”. This motion has an = 2.2 M yr, being in the range of the we derive for our VGs (see Table D.1).
@@
In Rawat et al. (2024) they analyze CO(10) ALMA data observing a V-shaped VG (see their Fig. 14) along the ridge of the G148.24+00.41 (G148) cloud. This V-shape peaks towards the dense clump at the center of this region, possibly indicating gas inflow along their filaments F2 and F6 towards the hub. Note that the length of the V-shape in G148 is pc, while our most prominent V-shape (Fig. 10) span and lengths of their identified filaments are 1.3 to 6.9 M and 14 to 38 pc respectively, large compared to the total mass (2.5 M) and extent of G353 ( pc). This different in probed length and masses might be reflected by their mean VG km s pc, orders of magnitude smaller than our VGs.
@@
In Pan et al. (2024) using APEX CO (21) data of the filamentary cloud G034.43+00.24 (G34) they identify converging VGs of lengths pc towards the “middle ridge” see their Fig. 3, top panel). They interpret these VG as gas flowing from the filaments onto dense clumps, located at the center of G34. These VGs of their southern and northern filaments are in the range of km s pc, and they estimate the total mass inflow rate towards the middle ridge as M yr, similar to our estimates.
@@
As presented here, V-shaped VGs appear to be a generic feature across a wide range of star forming environments, probing VGs with differences of up to orders of magnitude in spatial scales ranging from 0.1 to 10 pc. Based on these works, the V-shaped VGs appear to be a generic feature in star forming regions. Despite being commonly detected in recent studies, it is still not clear how they are produced. Henshaw et al. (2014) highlights the degeneracy regarding the opposite interpretations of these V-shaped velocity structures. They suggest that these VGs can be a signature of gas flows along kinked filaments towards a core located at their convergence point. From our analysis regarding the most prominent V-shape (see § 5.3 & 6) we see that no core is located at its apex, although cores 2 & 3 are within pc. Also the position difference between the center of the V-shape with the barycenter between these two cores is AU. This is consistent with the idea of small-scale gravitational collapse within the protocluster, similar to clump decoupling from their parent molecular cloud (Peretto et al., 2023). Based on this, we conclude that cores may be located in the vicinity of the velocity apex, and not necessarily on top of it. These gas flows towards denser regions may result in the formation of high-mass cores in later stages during the evolution of the protocluster.
@@
A different interpretation is that they are the product of out-flowing material coming from a forming protostar interacting with the surrounding dense gas (see their Fig. 12). To shed light into this degeneracy in G353 we compare the NH (dense gas tracer) radial velocities and SiO (shock/outflow tracer) as a proxy for energies. The velocity range V covered by the NH emission is 8 km s, while for SiO is 80 km s. Given the difference in probed velocities between SiO and NH (see § 5.3) and the analysis presented in § 6 we suggest that the V-shapes in G353 are a signature of infall.
@@
As presented in this section, the V-shaped VGs appear to be common in recent studies of protoclusters. The multiple VGs that conform the V-shapes present in G353 have values of to km s pc, with timescales ranging from to kyr, and values . These values are similar to VGs in other regions with comparable sizes and masses. In G353 it is likely that these VGs are the result of dense gas moving through filaments, possibly increasing the density of the central regions, shaping the overall velocity field at large and small scales, and leading to a further increase of the core population and star formation activity.
@@
One important aspect of the V-shapes that is still not well understood is the timescale associated of the VGs (VG). It is not clear if nor how these timescales determine core formation lifetimes or impact the star formation environment in general. In our sample of V-shapes the timescales are in tens of kyrs with the average value of kyr, times the cores , while the of the whole protocluster is Myr. In Rawat et al. (2024) they estimate the longitudinal collapse timescales for their filaments, being in the range of 5 15 Myr. Using their derived VGs we estimate their associated timescales to be between and 50 Myr, orders of magnitude larger than our small scale V-shapes timescales. We suggest that the VG timescales might serve as an upper limit for filamentary collapse timescales. In Zhou et al. (2022) they determine gas accretion times as a function of the lengths of their filaments, assuming that the VGs produced by gas inflow (see their Fig. 11). At filament lengths comparable to our V-shapes ( pc) their gas accretion timescales are on the order of our estimates (see Table D.1).
@@
It is also interesting to consider the mass accretion rates measured here compared to the available protocluster mass reservoir to explore implications for the duration of the gas dominated phase. The total mass accretion rate of our V-shaped structures is 310 M yr (see Table D.1). Considering the total mass (M) of G353 as a mass reservoir, we estimate the time needed fully consume the gas (). Here we assume that is representative of flows feeding gas onto cores. We estimate M, where the total mass of the region is 2500 M (Motte et al., 2022). We obtain a = 0.8 Myr, of similar magnitude but about four times larger than the of the protocluster. Considering that our estimate of is certainly a lower limit (see discussion above), the actual value value of is likely to be shorter, so closer to the free-fall time estimate. Given that the protocluster does not appear to be in a state of free-fall (see § 6) but instead undergoing comparatively slow gravitational contraction, the similarity in these relatively crude estimates seems remarkable. While we do not yet have an explanation for why this relatively good match in timescales, it would seem to indicate that protocluster evolution may be a self-regulating process. Larger samples and similar analysis will test this hypothesis.
@@ 9 Summary & conclusions
We characterize the complex dense gas kinematics of G353 using ALMA-IMF LP observations. The data used in this paper mainly consist of the fully combined NH data cube, but we also include 1.3 mm continuum cores and DCN cores velocity catalogues, SiO 12 m observation, and a (H) 1.3 mm continuum derived map. We summarize our main results below.
@@
With our NH isolated component modeling, we find that most of the 1.3 mm cores are located in regions with 2 to 3 velocity components. This indicates kinematic complexity down to 4 kau scales.
@@
We increase the number of cores with a V estimate in this region by further examining the DCN emission and comparing it with the NH data extracted towards the core positions. We find that 11 cores, which were previously undetected in the in the DCN background-subtracted fitting from Cunningham et al. (2023), are identified with our method. With this approach we increase our core velocities sample from 15 to 26, accounting for of the total 45 1.3 mm continuum cores. These are presented in Table C.1.
@@
From the PV diagrams, we see the DCN core velocities are in agreement with the NH velocity distribution (within a few DCN channel widths). This suggests coupling between the cores and the dense gas in which they formed.
@@
In the intensity-weighted PVs we see clear V-shaped velocity structures, composed by two linear velocity gradients (VGs) converging into a common point. These VGs are present across all NH velocity components. Some of them are near the location of cores in both position and velocity (see § 5.2)
@@
We successfully characterize nine V-shaped VGs well detected in our NH data (see Fig. D.2).
@@
We estimate the barycenter of cores 2 & 3, presenting an offset relative to the center of the V-shape of ″( 600 au) well below the beam size of our NH data.
@@
For V-shape “B” we find that core “7”, with a mass of 6 M, is located within a beam size from its apex.
@@
We characterize the VGs composing our sample of V-shapes by applying linear fits to these distributions. We estimate timescales associated to the VGs as VG. These timescales are between 35 to 170 kyr, with an average of 67 kyr. These values are short compared to the t of the protocluster ( Myr), and times larger that the cores average t ( kyr).
@@
We suggest that at small scales the NH V-shaped structures indicate gas motions along filaments, towards denser regions.
@@
Using an H mass map and the V-shapes mean timescales, we derive H mass accretion rates of , consistent with previous studies on regions that present gas flows along filaments towards denser object or regions, such as protostars and clumps.
@@
In SiO, the PV structure covers a velocity range (V) of km s, while for NH V is km s. This difference suggests that NH is tracing infall, a less energetic processes compared to SiO, a shock and outflow tracer.
@@
We model a gravitationally collapsing sphere. The derived radial velocities are consistent with the large scale morphology of the traditional PV diagram. This agreement suggests that at large scales the G353 protocluster is undergoing gravitational contraction.

### v2

@@ 7 Mass accretion rates in the V-shaped structure
Based on the idea that the V-shapes are a result of gas flowing toward cores, in this section we provide estimates of their mass accretion rates () for N2H+ and H2.
7.1 H2 mass accretion rate
@@
To ensure that we estimate the V-shape (H2) on the same area as in the N2H+ hyperfine line fitting, we use the CASA task imregrid to obtain the continuum-derived (H2) map from Díaz-González et al. (2023) at the resolution of the N2H+ data. We determine that in this V-shape the total (H 1.17 cm-2 in an area of 0.013 pc2. Here, we derive a (H2) map using Eq. 5:
@@
where from this (H2) map we consider only the points that are part of the V-shape. To determine the mass associated to flowing motions we subtract the core masses (from Louvet et al. submitted) that are located inside this V-shape. Note here that this mass map is an upper limit given that we do not apply a background correction. We obtain a total of (H 53 M⊙. Considering used in Eq. 8, we derive the (H2) as:
We use the procedure described in this section to estimate the (H2) of other eight V-shapes shown in Fig. D.2. We include these values in Table D.1. The average (H2) of these V-shapes is 3.4 10-4 M⊙ yr-1. Note that V-shapes “H”, “C”, “F”, and “B” present the largest (H2) and they are located near or at the convergence point of the filaments (see Fig. 13)
@@
For comparison, using the core masses from Louvet et al. (submitted) we estimate the free-fall time of all 45 1.3 mm cores and their mass accretion rates. These values present large scattering, ranging from M⊙ yr-1, with 28 of these cores presenting (H2) M⊙ yr-1. For cores 2 & 3, the average (H2) is 15.5 M⊙ yr-1, about twice the (H2) of the main blue V-shape, located between these two cores.
7.2 N2H+ mass accretion rate and relative abundance
To derive the associated to the main blue V-shape (Fig. 10), we need to estimate its N2H+ mass. For this procedure we use PySpecKit with the n2hp_vtau fitter, to fit the full N2H+ line. The fitted parameters (see below) allow us to derive the (N2H+). The V-shape structure contains 77, 143, and 25 N2H+ spectra with one, two, and three velocity components respectively. The bluest velocities in the three velocity component spectra accounts for 2% of the total number of velocities in this V-shape. For this reason, we model the full N2H+ hyperfine line structure with one and two velocity components. In Table 2 we list the parameters and ranges used for this procedure.
After obtaining the modeled N2H+ cube, we remove modeled components where , where and represent the estimated opacity and its associated error, respectively. This criterion is to ensure that we use reliable fitted parameters to determine our (N2H+) values. For the fitting of two N2H+ components we only analyze the most blueshifted component. The resulting opacities follow a log-normal distribution with a peak at .
We derive the (N2H+) of the V-shape by using Eq. 7 (Caselli et al., 2002b):
Where , Tex, and are the opacity, excitation temperature, and velocity dispersion respectively, obtained from the full line fitting. The Planck and Boltzmann constants are represented by and respectively, and are the frequency and wavelength of N2H+, is the Einstein coefficient of the N2H+ (10) transition, and are the statistical weights of the lower and upper energy levels, is the energy of the lower level, is the partition function estimated using the excitation temperature of the full N2H+ fits (see Eq. A2 from Caselli et al., 2002b).
From the above procedure, inside the V-shaped structure, we get a total (N2H+) = 5.241015 cm-2 and a total (N2H+) = 5.910-8 M⊙, within its extent of 0.013 pc2. We use the average timescale of the VGs from § 5.3 (Fig. 10), = 64.5 kyr, to determine the (N2H+) as:
Note that the (N2H+) estimate (and (H2) below) should be multiplied by , in order to account for the unknown inclination angle () of the protocluster/filaments relative to the POS.
@@ 7.2 N2H+ mass accretion rate and relative abundance
For the main blue V-shape (Fig. D), we derive the N2H+ relative abundance (N2H+), using the N2H+ and H2 column densities obtained above, as:
The N2H+) value obtained above appears lower than typical estimates in massive Galactic star forming regions reported in different works (; Caselli et al., 2002a; Henshaw et al., 2014, Sandoval-Garrido et al. in prep.). We consider it a good agreement considering the uncertainties of the involved measurements (i.e. column density estimates). For a comparison of the N2H+ relative abundance between regions composing the V-shapes and the rest of protocluster we would require to model the full N2H+ line emission in the whole region.
Excitation temperature [K] | (Tex) | 2.73 80 | |
Opacity | () | 0.01 40 | |
Centroid velocity [km s-1] | () | 25 15 | |
Velocity dispersion [km s-1] | () | 0.20 3 |
@@ 8.1 V-shaped VGs in the literature
The V-shaped velocity gradients described in this work have been detected across multiple Galactic star forming system. Stutz & Gould (2016) introduced the Slingshot mechanism in the Integral Shaped Filament (ISF) located in Orion A. They show undulations of the region in both position and velocity, suggesting that these features appear to be ejecting protostars (see their Fig. 2). Furthermore, Stutz (2018) characterize a standing wave in the neighborhood of the ISF, consistent with the Slingshot mechanism. It is possible that the undulations in the works above might result in the observed V-shaped structures seen in different studies (see below). González Lobos & Stutz (2019) identify six evenly spaced (every 0.44 pc) velocity peaks along the spine of the ISF in Orion A. They suggest that this periodicity is consistent with the wave-like perturbation in the gas caused by the Slingshot mechanism. In Álvarez-Gutiérrez et al. (2021) they analyze the L1482 filament located in the California Molecular Cloud. In all of the analyzed tracers there is a clear velocity peak in their north region (length 1.8 pc, mass M⊙). This subregion contains a higher gas density and higher number of YSOs compared to the more quiescent south part.
@@
While the two regions described above are considered nearby (both at D 500 pc), more distant regions also present these velocity features which we list below. Zhou et al. (2022) study the velocity profiles along filaments from the ATOMS survey (Liu et al., 2020b). The median mass of their sources is 1.4 M⊙ with a median length of the filaments at pc. By analyzing the H13CO+ (10) emission they find converging VGs along filaments (see their Figs. 6 & 10), which they also detected using simulations from Gómez & Vázquez-Semadeni (2014). These VGs at scales comparables to the V-shapes presented here are consistent with our VGs estimates (see their Fig. 7 & 8). In Zhou et al. (2023) they analyzed 13CO (21) APEX/LAsMA data of the G333 complex. They identify multiple V-shaped VG (see their Fig. 7) which they describe as the PV projection of a funneling structure in PPV space (see their Fig. 9). The origin of this structure is due to material inflowing towards the central hub and also due to gravitational contraction of star-forming clouds or clumps.
@@
Redaelli et al. (2022) use ALMA N2H+ (10) isolated component data of the high-mass (5200 M⊙) clump AGAL014.492-00.139 identifying multiple coherent structures in PPV space (trees “B” and “G”; right panel of their Fig. 7 & 9). These are characterized by multiple undulations, and possible V-shaped VGs. For their “G” PV distribution, they suggest that one scenario is where the dense gas is flowing along the filament (of length 0.2 pc) from protostar “p3” towards the protostar “p2”. This motion has an = 2.2 M⊙ yr-1, being in the range of the we derive for our VGs (see Table D.1).
@@
In Rawat et al. (2024) they analyze 13CO(10) data obtained with the Purple Mountain Observatory, as part of the Milky Way Imaging Scroll Painting survey. They detect a V-shaped VG (see their Fig. 14) along the ridge of the G148.24+00.41 (G148) cloud. This V-shape peaks towards the dense clump at the center of this region, possibly indicating gas inflow along their filaments F2 and F6 towards the hub. Note that the length of the V-shape in G148 is pc, while our most prominent V-shape (Fig. 10) is . Also the masses and lengths of their identified filaments are (1.3 to 6.9) M⊙ and 14 to 38 pc respectively, large compared to the total mass (2.5 M⊙) and extent of G353 ( pc). This difference in probed lengths and masses might be reflected by their mean VG km s-1 pc-1, orders of magnitude smaller than our VGs. This is consistent with the analysis presented in Zhou et al. (2022, 2023), where they observe an inverse relation between the spatial scale of a region and their velocity gradients (see their Fig. 8).
@@
In Pan et al. (2024) using APEX C18O (21) data of the filamentary cloud G034.43+00.24 (G34) they identify converging VGs of lengths pc towards the “middle ridge” see their Fig. 3, top panel). They interpret these VG as gas flowing from the filaments onto dense clumps, located at the center of G34. These VGs of their southern and northern filaments are in the range of km s-1 pc-1, and they estimate the total mass inflow rate towards the middle ridge as M⊙ yr-1, similar to our estimates.
@@
8.2 Filamentary 3D morphology
V-shaped VGs appear to be a generic feature across a wide range of star forming environments, probing VGs with differences of up to orders of magnitude in spatial scales ranging from 0.1 to 10 pc. Despite being commonly detected in recent studies, it is still not clear how they are produced. Henshaw et al. (2014) highlights the degeneracy regarding the opposite interpretations of these V-shaped velocity structures. They suggest that these VGs can be a signature of gas flows along kinked filaments towards a core located at their convergence point. From our analysis regarding the most prominent V-shape (see § 5.3 & 6) we see that no core is located at its apex, although cores 2 & 3 are within pc. Also the spatial offset between the center of the V-shape with the barycenter between these two cores is AU. This is consistent with the idea of small-scale gravitational collapse within the protocluster, similar to clump decoupling from their parent molecular cloud (Peretto et al., 2023). Based on this, we conclude that cores may be located in the vicinity of the velocity apex, and not necessarily on top of it. These gas flows towards denser regions may result in the formation of high-mass cores in later stages during the evolution of the protocluster.
@@ 8.2 Filamentary 3D morphology
A different interpretation is that they are the product of out-flowing material coming from a forming protostar interacting with the surrounding dense gas (see their Fig. 12). To shed light into this degeneracy in G353 we compare the N2H+ (dense gas tracer) radial velocities and SiO (shock/outflow tracer) as a proxy for energies. The velocity range V covered by the N2H+ emission is 8 km s-1, while for SiO is 80 km s-1. Given the difference in probed velocities between SiO and N2H+ (see § 5.3) and the analysis presented in § 6 we suggest that the V-shapes in G353 are a signature of infall.
@@
The multiple VGs that conform the V-shapes present in G353 have values of to km s-1 pc-1, with timescales ranging from to kyr, and values . These values are similar to VGs in other regions with comparable sizes and masses. In G353 it is likely that these VGs are the result of dense gas moving through filaments, possibly increasing the density of the central regions, shaping the overall velocity field at large and small scales, and leading to a further increase of the core population and star formation activity.
8.3 Timescales and mass accretion rates
@@
One important aspect of the V-shapes that is still not well understood is the timescale associated of the VGs (VG-1). It is not clear if nor how these timescales determine core formation lifetimes or impact the star formation environment in general. In our sample of V-shapes the timescales are in tens of kyrs with the average value of kyr, times the cores , while the of the whole protocluster is Myr. In Rawat et al. (2024) they estimate the longitudinal collapse timescales for their filaments, being in the range of 5 15 Myr. Using their derived VGs we estimate their associated timescales to be between and 50 Myr, orders of magnitude larger than our small scale V-shapes timescales. We suggest that the VG timescales might serve as an upper limit for filamentary collapse timescales. In Zhou et al. (2022) they determine gas accretion times as a function of the lengths of their filaments, assuming that the VGs produced by gas inflow (see their Fig. 11). At filament lengths comparable to our V-shapes ( pc) their gas accretion timescales are on the order of our estimates (see Table D.1).
8.4 Depletion timescales
@@
It is also interesting to consider the mass accretion rates measured here compared to the available protocluster mass reservoir to explore implications for the duration of the gas dominated phase. The total mass accretion rate of our V-shaped structures is 310-3 M⊙ yr-1 (see Table D.1). Considering the total mass (MTot) of G353 as a mass reservoir, we estimate the depletion timescale () as the time needed fully consume the gas. Here we assume that is representative of flows feeding gas onto cores. We estimate M, where the total mass of the region is 2500 M⊙ (Motte et al., 2022). We obtain a = 0.8 Myr, of similar magnitude but about four times larger than the of the protocluster. Considering that our estimate of is certainly a lower limit (see discussion above), the actual value value of is likely to be shorter, so closer to the free-fall time estimate. Given that the protocluster does not appear to be in a state of free-fall (see § 6) but instead undergoing comparatively slow gravitational contraction, the similarity in these relatively crude estimates seems remarkable. While we do not yet have an explanation for why relatively good match in timescales, it would seem to indicate that protocluster evolution may be a self-regulating process. Larger samples and similar analysis will test this hypothesis.
@@ 9 Summary & conclusions
We characterize the complex dense gas kinematics of G353 using ALMA-IMF LP observations. The data used in this paper mainly consist of the fully combined N2H+ data cube, but we also include 1.3 mm continuum cores and DCN cores velocity catalogues, SiO 12 m observation, and a (H2) 1.3 mm continuum derived map. We summarize our main results below.
@@
With our N2H+ isolated component modeling, we find that most of the 1.3 mm cores are located in regions with 2 to 3 velocity components. This indicates kinematic complexity down to 4 kau scales.
@@
We increase the number of cores with a VLSR estimate in this region by further examining the DCN emission and comparing it with the N2H+ data extracted towards the core positions. We find that 11 cores, which were previously undetected in the in the DCN background-subtracted fitting from Cunningham et al. (2023), are identified with our method. With this approach we increase our core velocities sample from 15 to 26, accounting for of the total 45 1.3 mm continuum cores. These are presented in Table C.1.
@@
From the PV diagrams, we see the DCN core velocities are in agreement with the N2H+ velocity distribution (within a few DCN channel widths). This suggests coupling between the cores and the dense gas in which they formed.
@@
In the intensity-weighted PVs we see clear V-shaped velocity structures, composed by two linear velocity gradients (VGs) converging into a common point. These VGs are present across all N2H+ velocity components. Some of them are near the location of cores in both position and velocity (see § 5.2)
@@
We estimate the barycenter of cores 2 & 3, presenting an offset relative to the center of the V-shape of ″( 600 au) well below the beam size of our N2H+ data.
@@
For V-shape “B” we find that core “7”, with a mass of 6 M⊙, is located within a beam size from its apex.
@@
We characterize the VGs composing our sample of V-shapes by applying linear fits to these distributions. We estimate timescales associated to the VGs as VG-1. These timescales are between 35 to 170 kyr, with an average of 67 kyr. These values are short compared to the tff of the protocluster ( Myr), and times larger that the cores average tff ( kyr).
@@
We suggest that at small scales the N2H+ V-shaped structures indicate gas motions along filaments, towards denser regions. Thus we interpret as inflow timescales.
@@
Using an H2 mass map and the V-shapes mean timescales, we derive H2 mass accretion rates of , consistent with previous studies on regions that present gas flows along filaments towards denser object or regions, such as protostars and clumps. Moreover, V-shapes “H”, “C”, “F”, and “B” present the largest (H2) and they are located near or at the convergence point of the filaments (see Fig. 13).
@@
In SiO, the PV structure covers a velocity range (V) of km s-1, while for N2H+ V is km s-1. This difference suggests that N2H+ is tracing infall, a less energetic processes compared to SiO, a shock and outflow tracer.
@@
We model the protocluster as a gravitationally collapsing sphere. The derived radial velocities are consistent with the large scale morphology of the traditional PV diagram. This agreement suggests that at large scales the G353 protocluster is undergoing gravitational contraction.